# A TextGCN-Based Decoding Approach for Improving Remote Sensing Image Captioning

### v2

@@ Abstract
Remote sensing images are highly valued for their ability to address complex real-world issues such as risk management, security, and meteorology. However, manually captioning these images is challenging and requires specialized knowledge across various domains. This letter presents an approach for automatically describing (captioning) remote sensing images. We propose a novel encoder-decoder setup that deploys a Text Graph Convolutional Network (TextGCN) and multi-layer LSTMs. The embeddings generated by TextGCN enhance the decoder’s understanding by capturing the semantic relationships among words at both the sentence and corpus levels. Furthermore, we advance our approach with a comparison-based beam search method to ensure fairness in the search strategy for generating the final caption. We present an extensive evaluation of our approach against various other state-of-the-art encoder-decoder frameworks. We evaluated our method across three datasets using seven metrics: BLEU-1 to BLEU-4, METEOR, ROUGE-L, and CIDEr. The results demonstrate that our approach significantly outperforms other state-of-the-art encoder-decoder methods.
@@ II-C Text Graph Convolution Network
Here, the adjacency matrix A is constructed based on the Point-wise Mutual Information (PMI) between two words, using a fixed-size window applied to each sentence. The degree matrix is denoted as D which is a diagonal matrix calculated by counting non zero elements in the corresponding row of the adjacency matrix. $F_{i}$ represents the feature matrix and $W^{i}$ denotes the weight matrix at the $i^{th}$ layer. The feature matrix can be initialized in various ways, with one popular method being the initialization of a diagonal matrix where the diagonal elements are the sum of the corresponding rows. However, since our aim is to find the embedding matrix, we have applied some extensions. We take the ceiling value of the diagonals and then project it onto an embedding initializer (torch.nn.Embedding in our work). The weight matrices $W_{0}$ and $W_{1}$ are initialized using the Normalized Xavier [22] weight initialization. Both layers apply a sigmoid activation function and standard Gaussian normalization to each row. At the end of this process, each row is divided by its maximum absolute value. The resulting feature matrix values from TextGCN serve as pre-calculated weights for the embedding layer, which is non-trainable (refer to Figure 1).
@@ II-E Multi-Layer Decoding Strategy
A conventional encoder-decoder model typically uses a single decoder that takes the embedding matrix of the text input sequence and produces an output text feature vector. This output text feature vector is then concatenated (or added) with the input image feature vector. A stack of linear layers processes it to predict the next word. We have extended this idea by incorporating two LSTMs at different stages (refer to Fig. 1). The concept adhered is that, instead of directly sending the concatenated output of the first LSTM and the image feature vectors to the stack of linear layers (SL), we pass it through another LSTM. This multi-layer decoding strategy [24] helps the model gain a better understanding of the image, as its output is influenced by both the image and corresponding text features. Finally, the SL layer handles the decision-making process.
@@ III-A Dataset and Performance Metrics
In this study, we utilized three publicly available remote sensing image dataset: RSICD [25]UCM [26], and SYDNEY [26]. Our analysis of these dataset revealed spelling and grammatical errors, as well as diverse country-specific variations in certain words within the dataset. To ensure consistency, we corrected these issues by standardizing to American English conventions. The updated dataset are also referenced in the supplementary file. Our captioning system is thoroughly evaluated across three dataset using seven metrics: BLEU-1 to BLEU-4, METEOR, ROUGE-L, and CIDEr. The evaluation involves calculating BLEU, METEOR, and CIDEr with the nlgeval [27] package, while the ROUGE-L metric is computed using standard Python modules. In this letter, we have presented results exclusively from the RSICD dataset.
@@ III-C1 The Performance of the Proposed Method
TABLE I: Experimental Results for Captioning on RSICD Dataset (M-T-C is our proposed method)
@@
Table I compares various baseline models and search techniques on the RSICD dataset. In the first section of Table I , we present the results for 1. RNN with Bag of Words features representation (RBOW) [25], 2. LSTM with Fisher Vector features representation (LFV) [25], 3. Collective Semantic Metric Learning Framework (CSMLF) [29], 4. Merged GRU Decoder model (MGRUD) [8], and 5. SVM based Decoder model (SVMDC) [8]. The subsequent sections evaluate different search techniques. The naming convention used is as follows: model (E for conventional encoder-decoder and M for the multi-layer decoder-based RSIC model used in our work) - embeddings (N for no pretrained embeddings and T for TextGCN embeddings) - search method ( B for traditional beam search, and C for the comparison-based beam search [17] after modifications). Table I shows that the combination of the multi-layer decoder, TextGCN embeddings, and comparison-based beam search (M-T-C) produces the best results. We observed that the multi-layer decoder performs better than the conventional encoder-decoder models, and the comparison-based beam search also improves the quality of the captions compared to traditional searching techniques.

### v3

@@ Abstract
Remote sensing images are highly valued for their ability to address complex real-world issues such as risk management, security, and meteorology. However, manually captioning these images is challenging and requires specialized knowledge across various domains. This letter presents an approach for automatically describing (captioning) remote sensing images. We propose a novel encoder-decoder setup that deploys a Text Graph Convolutional Network (TextGCN) and multi-layer LSTMs. The embeddings generated by TextGCN enhance the decoder’s understanding by capturing the semantic relationships among words at both the sentence and corpus levels. Furthermore, we advance our approach with a comparison-based beam search method to ensure fairness in the search strategy for generating the final caption. We present an extensive evaluation of our approach against various other state-of-the-art encoder-decoder frameworks. We evaluated our method using seven metrics: BLEU-1 to BLEU-4, METEOR, ROUGE-L, and CIDEr. The results demonstrate that our approach significantly outperforms other state-of-the-art encoder-decoder methods.
@@ II-C Text Graph Convolution Network
Here, the adjacency matrix A is constructed based on the Point-wise Mutual Information (PMI) between two words, using a fixed-size window applied to each sentence. The degree matrix is denoted as D which is a diagonal matrix calculated by counting non zero elements in the corresponding row of the adjacency matrix. $F_{i}$ represents the feature matrix and $W^{i}$ denotes the weight matrix at the $i^{th}$ layer. The feature matrix can be initialized in various ways, with one popular method being the initialization of a diagonal matrix where the diagonal elements are the sum of the corresponding rows. However, since our aim is to find the embedding matrix, we have applied some extensions. We have taken the ceiling value of the diagonals and then project it onto an embedding initializer (torch.nn.Embedding in our work). The weight matrices $W_{0}$ and $W_{1}$ are initialized using the Normalized Xavier [22] weight initialization. Both layers apply a sigmoid activation function and standard Gaussian normalization to each row. At the end of this process, each row is divided by its maximum absolute value. The resulting feature matrix values from TextGCN serve as pre-calculated weights for the embedding layer, which is non-trainable (refer to Figure 1).
@@ II-E Multi-Layer Decoding Strategy
A conventional encoder-decoder model typically uses a single decoder that takes the embedding matrix of the text input sequence and produces an output text feature vector. This output text feature vector is then concatenated (or added) with the input image feature vector. A stack of linear layers processes it to predict the next word. We have extended this idea by incorporating two LSTMs at different stages (refer to Fig. 1). The concept adhered is that, instead of directly sending the concatenated output of the first LSTM and the image feature vectors to the stack of linear layers (SL), it is passed through another LSTM. This multi-layer decoding strategy [24] helps the model gain a better understanding of the image, as its output is influenced by both the image and corresponding text features. Finally, the SL layer handles the decision-making process.
@@ III-A Dataset and Performance Metrics
In this study, we used three publicly available remote sensing image dataset: RSICD [25]UCM [26], and SYDNEY [26]. RSICD is a balanced dataset with a reasonably good number of instances in comparison to the other two dataset [25]. Hence, in this letter, we have presented results exclusively from the RSICD dataset. Our analysis of these dataset revealed spelling and grammatical errors, as well as diverse country-specific variations in certain words within the dataset. To ensure consistency, we have corrected these issues by standardizing to American English conventions. The updated dataset are also referenced in the supplementary file. Our captioning system is thoroughly evaluated using seven metrics: BLEU-1 to BLEU-4, METEOR, ROUGE-L, and CIDEr. The evaluation involves calculating BLEU, METEOR, and CIDEr with the nlgeval [27] package, while the ROUGE-L metric is computed using standard Python modules.
@@ III-C1 The Performance of the Proposed Method
TABLE I: Experimental Results for Captioning on RSICD Dataset
@@
Table I compares various baseline models and search techniques on the RSICD dataset. In the first section of Table I , we present the results for 1. RNN with Bag of Words features representation (RBOW) [25], 2. LSTM with Fisher Vector features representation (LFV) [25], 3. Collective Semantic Metric Learning Framework (CSMLF) [29], 4. Merged GRU Decoder model (MGRUD) [8], and 5. SVM based Decoder model (SVMDC) [8]. The subsequent sections evaluate different search techniques. The naming convention used is as follows: model (E for conventional encoder-decoder and M for the multi-layer decoder-based RSIC model used in our work) - embeddings (N for no pretrained embeddings and T for TextGCN embeddings) - search method ( B for traditional beam search, and C for the comparison-based beam search [17] after modifications). Table I demonstrates that the proposed method, which combines the multi-layer decoder, TextGCN embeddings, and comparison-based beam search (M-T-C), yields the best results. We observed that the multi-layer decoder performs better than the conventional encoder-decoder models, and the comparison-based beam search also improves the quality of the captions compared to traditional searching techniques.